OpenStudy (anonymous):

Your cap size is based on the circumference of your head (in inches). For a certain range of head circumferences c, cap size s can be modeled by the following equation: s = c - 1/3. If you wear a size 6.5 cap what is your head circumferences?

OpenStudy (jlastino):

You wear a 6.5 cap, meaning "s" would be equal to 6.5 6.5=c-(1/3) Multiply both sides by3 3(6.5)= c(3)- 1/3(3) 19.5=3c-1 19.5+1=3c 20.5=3c c=6.83